You are considering investing $1000 in a complete portfolio. The complete portfolio is composed of Treasury notes that pay 5% and a risky portfolio, P, constructed with two risky securities X and Y. The optimal weights of X and Y in P are 60% and 40% respectively. X has an expected rate of return of 14% and Y has an expected rate of return of 10%. If you decide to hold 25% of your complete portfolio in the risky portfolio and 75% in the Treasury notes then calculate the dollar values of your positions in X and Y respectively?
Answer #1

Answer - Suppose Mr X is considering investment of portfolio

We have 3 asset i.e. treasury , X stock and Y stock

Optimal weights of portfolio

Treasury = 75%

Rest Mr x wants to put 25% in X and Y

Optimal weights of risky securities

X = 25% *0.6

( as mentioned in the question about optimal weights)

= 15%

Y = 25%*0.4

= 10%

Thus treasury = 75%

X Stock = 15%

Y Stock = 10%

If Mr. X has $1000 to invest, then proportionate amount is

Treasury = 1000*75%

= $750

X Stock = 1000*15%

= $150

Y stock = 1000*10%

=$100

Thus dollar value of X and Y are $150 and $100 respectively.

expected return = 0.75 * treasury + 0.15*X stock + 0.10*Ystock

= 0.75(5%)+ 0.15(14%) +0.10(10%)

=6.85%